CARA Telecon Minutes Date: April 2, 2001 Time: 1:00 PM Attendance: BU: T. Bania Cornell: G. Stacey CMU: J. Peterson NU: G. Novak R.I.T: S.A.O.: T. Stark, W. Walsh South Pole: C. Martin, C. Kaminski Chicago: J. Carlstrom, A. Harper, S. Meyer, N. Odalen, R. Pernic, R. Landsberg, R. Loewenstein, J. Bausch, E. Leitch UCB: UCSB: J. Ruhl Approval of previous minutes: Motion made and seconded. Agenda Items News from the Pole: C. Martin reported that the wind speed has been over 20 knots. AST/RO shut down its observing during this period. AST/RO is now back to observing SMC in CO 2-1 The Viper Telescope is broken and M. Newcomb is looking into the problem. The AZ servo amp (Techron) has been replaced and this is not part of the current problem. J. Carlstrom commented that the amp was not to be replaced until all the problems of the original failure were know and reported. C. Kaminski stated that M. Newcomb had put on the protection circuits so the amp should not fail. The amp has protective switches that were not turned on the original unit. This new amp has all the switched turned on. M. Newcomb had looked over the failed amp and found that the amp was not properly set up. *** J. Carlstrom requested that the problems of the failure be written up and that the solutions be document. J. Ruhl and J. Peterson were requested to contact the team and as a team prepare in full detail a report that explains the problems that were found and the solutions. Acbar - has been off the air due to the debugging problems of the Viper Telescope. AST/RO - operating as weather permits. DASI - beam maps being observed; `tracks' taken to evaluate the effectiveness of ground shield for polarization observations are being done. A new version of the pointing model has been developed to correctly implement tilt correction at South Pole (old code does not work at Pole since it adjust effect longitude and lattitude of observatory). DASI should be taking target data full time shortly. SPIFI - Instrument is at the JCMT now waiting for telescope time on Friday. Everything fine. Hoping for good weather. SPARO - the instrument was worked on over the summer at Pole and this work has improved the instrument by a factor of 3 due to improved optical coupling. T. Bania reported that M. Huang has received his Ph.D. AIRO is in limbo waiting for word from NSF although grapevine says the reviews went well. No word on new proposals -- there are a total of six! R, Landsberg reminded the group that on April 6th the responses are due on the REU's. Current status on Liquid Helium - On 1 Wessington the loss rate is going down to 17 liters per day. The 3000-gallon dewar is up to a loss rate of 22 liters per day. T. Stark, G. Novak & J. Ruhl to get together and discuss helium issues. J. Ruhl will send around Art Francis's memo and "cliff notes." Equipment - J. Peterson has sent around the list of equipment that has been requested by CARA. There was a question regarding Techrons - should they be a part of the equipment list for each instrument? It was decided that they were part of each instrument. It was discussed and an agreement was made that CARA should request Raytheon to purchase a high end oscilloscopes and 2 mid-rage portable units with battery packs at a cost of approx. $5,000.00 each. *** S. Meyer to provide to J. Peterson specs on these oscilloscopes. Question was asked if any equipment should be purchased for the machine shop. R. Pernic responded in the negative. N. Odalen reported that Chicago has not received funds and that ** she would be checking with NSF and report to the group on the status of these funds.